# Optimal Rates of Convergence for Entropy Regularization in Discounted Markov Decision Processes

### v3

@@ 4.2. Convergence of policies
of the policy $\pi$ to the set of optimal policies. By concavity we have $-\log(1-h)\leq\frac{h}{1-h}$ for $h<1$ and using Lemma 4.6 we estimate
@@
Similar to the upper bound, we use $-\log(1-h)\geq h$ for $h<1$ and Lemma 4.6 to estimate
@@ 4.3. Convergence of tabular unregularized natural policy gradients
Our continuous time analysis in Lemma 4.3 is inspired by the proof of exponential convergence of unregularized natural policy gradient methods given in [28]. Here, we provide an essentially sharp analysis of unregularized natural policy gradient methods in discrete time that holds for all iterations.
@@
where $Z_{k}(s)=\sum_{a\in\mathscr{A}}\pi_{k}(a|s)e^{\eta(1-\gamma)^{-1}A^{\pi_{k}}(s,a)}$, see [1]. We now characterize the optimal exponential convergence rate for natural policy gradient methods in the following setting.
@@
Similar upper and lower bounds in an asymptotic regime were given in [28, 33], respectively, where, our guarantees hold for all iterates. The proof of Theorem 4.9 is analoguous to the analysis in continuous time, and is built on the following the discrete-time analogon of Proposition 3.11.
@@
Proposition 4.10 is an extension of the convergence analysis in [1, Theorem 16] and we defer its proof to Appendix B. Similar to Lemma 4.3 we provide exponential bounds on the entries $\pi_{k}(a|s)$ of the policies. Our analysis is similar to the one in [28, Lemma 3.4], however, our bound holds for arbitrary initial policies, for all iterations, and comes with a lower bound matching up to a polynomial term.
@@
The statements follow just like in the proof of Theorem 4.2 where here we combine Lemma 4.4 and Lemma 4.7 with Lemma 4.11. ∎
@@ 5. The Case of General Convex Regularizers
Although entropy regularization is arguably the most classic regularization strategy in reinforcement learning, more general regularizers have been considered in a variety of works [42, 19, 29, 52]. It is the purpose of this section to extend our explicit analysis of the entropy regularization to general convex regularizers. We show that for general convex regularizers considered in reinforcement learning, a corresponding linear programming formulation with convex regularization can be derived. In particular, this allows us to extend the implicit bias results to this more general class of policy mirror descent methods. Finally, we discuss for which regularizers one obtains an explicit expression of the policy gradient, allowing for an explicit convergence analysis. For ease of notation, we use the short hand notation $\pi_{s}=\pi(\cdot|s)$ for a policy $\pi\in\Delta_{\mathscr{A}}^{\mathscr{S}}$ as well as $\nu_{s}=\nu(s,\cdot)$ for a vector $\mathbb{R}^{\mathscr{S}\times\mathscr{A}}$.
@@
The perspective of $\psi$ is given by $f(\mu,t)=t\psi(\frac{\mu}{t})$, which is convex, see [9, Subsection 3.2.6]. As $\mu\mapsto(\mu,\sum_{a\in\mathscr{A}}\mu(a))$ is linear, $\phi$ is convex as the composition of a linear and convex function.
@@
Just like in the case of Theorem 3.3 we obtain that
@@
If further $h^{\prime\prime}(x)=x^{-\sigma}$ for $\sigma\in(1,+\infty)$ we can proceed analogously to the case of entropy regularization. Here, we obtain, together with the sublinear convergence, the differential inequality
@@
$\psi(\mu)=\begin{cases}\sum_{a\in\mathscr{A}}\log\mu(a)\quad&\text{if }\sigma=2\\ \frac{1}{(2-\sigma)(1-\sigma)}\sum_{a\in\mathscr{A}}\mu(a)^{2-\sigma}\quad&\text{otherwise}\end{cases}$ (5.21)
and consider
@@
First, we show the global well-posedness of the gradient flows in this case. Here, note that it suffices to show the Legendre-type property $\lVert\nabla\Phi(\nu_{k})\rVert\to+\infty$ for $\nu_{k}\to\nu\in\partial\mathscrsfs{D}$, see [3, 37]. Fix $\nu_{k}\to\nu\in\partial\mathscrsfs{D}$ and denote the corresponding state distributions by $d_{k}\to d$. Then we have $\nu_{k}(s,a)\to 0$ for some $s\in\mathscr{S},a\in\mathscr{A}$ and by 3.2 we have $d(s)>0$. Hence, we have $\pi_{k}(a|s)\to 0$. Recall that we can express the regularizer as $\Phi(\nu)=\sum_{s\in\mathscr{S}}\phi(\nu_{k}(s,\cdot))$ with $\phi$ defined in (5.6) and hence it remains to show that $\lVert\nabla\phi(\eta(s,\cdot))\rVert\to+\infty$. Direct computation yields
@@ 6. Improved Iteration Complexity for Regularized Natural Policy Gradients
for entropy-regularized natural policy gradients with stepsize $\eta=\frac{\log k}{k}>0$, see [12]. See also [45] for an $O(\frac{1}{t})$ guarantee of entropy-regularized natural policy gradient flows with $\tau=\frac{1}{t}$. In contrast, unregularized natural policy gradient achieves an exponential convergence rate of $\tilde{O}(e^{-\Delta\eta k})$, see Theorem 4.9 and [28, 33]. We use our estimate on the regularization error and obtain the following improved guarantee for entropy-regularized natural policy gradients.
@@
Proposition 6.2 (Lipschitz-continuity of the reward [37]).
@@
First, note that $R^{\star}-R(\pi)=R^{\star}-R(\pi^{\star}_{\tau})+R(\pi^{\star}_{\tau})-R(\pi)$. Using the Lipschitz continuity of the reward from Proposition 6.2 as well as Pinsker’s and Jensen’s inequality we estimate
@@
Lemma 6.3 can by used in combination with any result bounding $D_{\operatorname{K}}(\pi^{\star}_{\tau},\pi_{k})$ for a policy optimization technique. If only bounds on $R_{\tau}(\pi^{\star}_{\tau})-R_{\tau}(\pi_{k})$ are available, one can also use the local bound $D_{\operatorname{K}}(\pi^{\star}_{\tau},\pi)\leq\omega\tau^{-1}(R_{\tau}(\pi^{\star}_{\tau})-R_{\tau}(\pi))$, which holds in a neighborhood of $\pi^{\star}_{\tau}$ that depends on $\omega\in(0,1)$, see [40, Lemma 29]. Here, we limit our discussion to entropy-regularized natural policy gradient methods, which are known to converge linearly.
Theorem 6.4 (Convergence of entropy-regularized NPG, [12]).
Consider natural policy gradient with a tabular softmax policy parametrization, a fixed regularization strength $\tau>0$ and stepsize $0<\eta\leq\frac{1-\gamma}{\tau}$ and denote the iterates of the natural policy gradient updates by $(\pi_{k})_{k\in\mathbb{N}}$. Then for $k\in\mathbb{N}$ it holds that
@@
Through a direct combination of Lemma 6.3, Theorem 4.2, and Theorem 6.4 we obtain (6.3). ∎
Recently, a sharp asymptotic analysis of entropy-regularized natural policy gradient methods has been conducted in [33] showing $R_{\tau}^{\star}-R(\pi_{k})=O((1+\eta\tau)^{-2k})$ compared to the $O((1-\eta\tau)^{k})$ convergence of Theorem 6.4. However, an analogous estimate on the convergence of the policies, as well as a control on the entry times after which the rate holds, is missing, which prevents us from using it in our analysis. By Theorem 4.9 unregularized natural policy gradients achieve $O(k^{\gamma c}e^{-\Delta\eta k})$ convergence. Hence, although Theorem 6.1 improves existing $O(\frac{\log k}{k})$ guarantees, the provided rate is still asymptotically slower compared to unregularized natural policy gradients. Where we consider a fixed regularization and step size, $O(\gamma^{k})$-convergence was established when exponentially decreasing the regularization and increasing the step size [32], which can also be achieved by unregularized policy mirror descent, where the rate $O(\gamma^{k})$ is known to be optimal [24]. For the small stepsize limit $\eta\to 0$ the updates of the regularized natural policy gradient scheme follow the regularized Kakade gradient flow recently studied in [27]. These results complement the discrete-time analysis and ensure that $D_{\operatorname{K}}(\pi^{\star}_{\tau},\pi_{t})\leq e^{-\tau t}D_{\operatorname{K}}(\pi^{\star}_{\tau},\pi_{0})$, see [27, Equation (61)]. An analog treatment to the discrete-time case yields an overall estimate of $O(e^{-\sqrt{{2^{-1}\Delta t}}})$ for the regularized flow with strength $\tau=\sqrt{2\Delta t^{-1}}$, compared to the existing $O(t^{-1})$ guarantee in [45].
@@ 7. Conclusions
In this work we provide an explicit characterization of the error introduced by entropy regularization in infinite-horizon discounted Markov decision processes with finite state and action spaces. We show that the solutions of the regularized optimization problems satisfy a gradient flow equation with respect to a Riemannian metric common in natural policy gradient methods. We use this to provide upper and lower bounds on the regularization error that decrease exponentially in the inverse regularization strength and match up to polynomial factors with an instance-dependent exponent. Thereby, we characterize the optimal exponential convergence rate for entropy regularization in discounted Markov decision processes. Further, this correspondence allows us to identify the limit of this gradient flow as the generalized maximum entropy optimal policy, thereby characterizing the implicit bias of this gradient flow, which corresponds to a time-continuous version of the natural policy gradient method. We use our improved error estimates to show that for entropy-regularized natural policy gradient methods, the total error decays exponentially in the square root of the number of iterations.
@@ A.1. Performance difference lemma
We recall an expression of the difference between the rewards of two policies in terms of the advantage function and the state-action distribution and for a proof we refer to [1, Lemma 2].
@@ Appendix B Sublinear Convergence of Unregularized Natural Policy Gradient
Here, we generalize the sublinear convergence analysis for unregularized natural policy gradients of [1, Section 5.3] to non-uniform initial policies $\pi_{0}$ and unnormalized reward $r\in\mathbb{R}^{\mathscr{S}\times\mathscr{A}}$.
@@
First, note that the natural policy gradient iteration is independent of the initial distribution $\mu\in\Delta_{\mathscr{S}}$. Using the performance difference Lemma A.1 as well as the definition of the natural policy gradient step we estimate the suboptimality for an arbitrary state $s_{0}\in\mathscr{S}$ via
@@ Acknowledgments
The authors acknowledge funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under the project number 442047500 through the Collaborative Research Center Sparsity and Singular Structures (SFB 1481). SC acknowledges funding by the Federal Ministry of Education and Research (BMBF) and the Ministry of Culture and Science of the German State of North Rhine-Westphalia (MKW) under the Excellence Strategy of the Federal Government and the Länder.

### v5

@@ 4.2. Convergence of policies
Consider 4.1 and set $c\coloneqq(1-\gamma)^{-1}D_{\operatorname{K}}(\pi^{\star},\pi_{0})$. For any $t\geq 1$, it holds that
@@
of the policy $\pi$ to the set of optimal policies. By concavity we have $\log(1-h)\leq\frac{h}{1-h}$ for $h<1$ and using Lemma 4.6 we estimate
@@
Similar to the upper bound, we use $\log(1-h)\geq h$ for $h<1$ and Lemma 4.6 to estimate
@@ 4.3. Convergence of tabular unregularized natural policy gradients
Our continuous time analysis in Lemma 4.3 is inspired by the proof of exponential convergence of unregularized natural policy gradient methods given in [khodadadian2022linear]. Here, we provide an essentially sharp analysis of unregularized natural policy gradient methods in discrete time that holds for all iterations.
@@
where $Z_{k}(s)=\sum_{a\in\mathscr{A}}\pi_{k}(a|s)e^{\eta(1-\gamma)^{-1}A^{\pi_{k}}(s,a)}$, see [agarwal2021theory]. We now characterize the optimal exponential convergence rate for natural policy gradient methods in the following setting.
@@
Similar upper and lower bounds in an asymptotic regime were given in [khodadadian2022linear, liu2024elementary], respectively, where our guarantees hold for all iterates. The proof of Theorem 4.9 is analogous to the analysis in continuous time, and is built on the following the discrete-time analog of Proposition 3.11.
@@
Proposition 4.10 is an extension of the convergence analysis in [agarwal2021theory, Theorem 16] and we defer its proof to Appendix B. Similar to Lemma 4.3, we provide exponential bounds on the entries $\pi_{k}(a|s)$ of the policies. Our analysis is similar to the one in [khodadadian2022linear, Lemma 3.4], however, our bound holds for arbitrary initial policies, for all iterations, and comes with a lower bound matching up to a polynomial term.
@@
The statements follow just like in the proof of Theorem 4.2, where here we combine Lemma 4.4 and Lemma 4.7 with Lemma 4.11. ∎
@@ 5. The Case of General Convex Regularizers
Although entropy regularization is arguably the most classic regularization strategy in reinforcement learning, more general regularizers have been considered in a variety of works [neu2017unified, geist2019theory, lan2022policy, zhan2021policy]. It is the purpose of this section to extend our explicit analysis of the entropy regularization to general convex regularizers. We show that for general convex regularizers considered in reinforcement learning, a corresponding linear programming formulation with convex regularization can be derived. In particular, this allows us to extend the implicit bias results to this more general class of policy mirror descent methods. Finally, we discuss for which regularizers one obtains an explicit expression of the policy gradient, allowing for an explicit convergence analysis. For ease of notation, we use the shorthand notation $\pi_{s}=\pi(\cdot|s)$ for a policy $\pi\in\Delta_{\mathscr{A}}^{\mathscr{S}}$ as well as $\nu_{s}=\nu(s,\cdot)$ for a vector $\mathbb{R}^{\mathscr{S}\times\mathscr{A}}$.
@@
The perspective of $\psi$ is given by $f(\mu,t)=t\psi(\frac{\mu}{t})$, which is convex, see [boyd2004convex, Subsection 3.2.6]. As $\mu\mapsto(\mu,\sum_{a\in\mathscr{A}}\mu(a))$ is linear, $\phi$ is convex as the composition of a linear and convex function.
@@
Just like in the case of Theorem 3.3, we obtain that
@@
If further $h^{\prime\prime}(x)=x^{-\sigma}$ for $\sigma\in(1,+\infty)$, we can proceed analogously to the case of entropy regularization. Here, we obtain, together with the sublinear convergence, the differential inequality
@@
$\psi(\mu)=\begin{cases}\sum_{a\in\mathscr{A}}\log\mu(a)\quad&\text{if }\sigma=2\\ \frac{1}{(2-\sigma)(1-\sigma)}\sum_{a\in\mathscr{A}}\mu(a)^{2-\sigma}\quad&\text{otherwise}.\end{cases}$ (5.21)
Further,
@@
First, we show the global well-posedness of the gradient flows in this case. Here, note that it suffices to show the Legendre-type property $\lVert\nabla\Phi(\nu_{k})\rVert\to+\infty$ for $\nu_{k}\to\nu\to\partial\mathscrsfs{D}$, see [alvarez2004hessian, mueller2023thesis]. Fix $\nu_{k}\to\nu\partial\mathscrsfs{D}$ and denote the corresponding state distributions by $d_{k}\to d$. Then we have $\nu_{k}(s,a)\to 0$ for some $s\in\mathscr{S},a\in\mathscr{A}$ and by 3.2 we have $d(s)>0$. Hence, we have $\pi_{k}(a|s)\to 0$. Recall that we can express the regularizer as $\Phi(\nu)=\sum_{s\in\mathscr{S}}\phi(\nu_{k}(s,\cdot))$ with $\phi$ defined in (5.6) and hence it remains to show that $\lVert\nabla\phi(\eta(s,\cdot))\rVert\to+\infty$. Direct computation yields
@@ 6. Improved Iteration Complexity for Regularized Natural Policy Gradients
for entropy-regularized natural policy gradients with stepsize $\eta=\frac{\log k}{k}>0$, see [cen2021fast]. See also [sethi2024entropy] for an $O(\frac{1}{t})$ guarantee of entropy-regularized natural policy gradient flows with $\tau=\frac{1}{t}$. In contrast, unregularized natural policy gradient achieves an exponential convergence rate of $\tilde{O}(e^{-\Delta\eta k})$, see Theorem 4.9 and [khodadadian2022linear, liu2024elementary]. We use our estimate on the regularization error and obtain the following improved guarantee for entropy-regularized natural policy gradients.
@@
Proposition 6.2 (Lipschitz-continuity of the reward [mueller2023thesis]).
@@
First, note that $R^{\star}-R(\pi)=R^{\star}-R(\pi^{\star}_{\tau})+R(\pi^{\star}_{\tau})-R(\pi)$. Using the Lipschitz continuity of the reward from Proposition 6.2 as well as Pinsker’s and Jensen’s inequality, we estimate
@@
Lemma 6.3 can by used in combination with any result bounding $D_{\operatorname{K}}(\pi^{\star}_{\tau},\pi_{k})$ for a policy optimization technique. If only bounds on $R_{\tau}(\pi^{\star}_{\tau})-R_{\tau}(\pi_{k})$ are available, one can also use the local bound $D_{\operatorname{K}}(\pi^{\star}_{\tau},\pi)\leq\omega\tau^{-1}(R_{\tau}(\pi^{\star}_{\tau})-R_{\tau}(\pi))$, which holds in a neighborhood of $\pi^{\star}_{\tau}$ that depends on $\omega\in(0,1)$, see [muller2024geometry, Lemma 29]. Here, we limit our discussion to entropy-regularized natural policy gradient methods, which are known to converge linearly.
Theorem 6.4 (Convergence of entropy-regularized NPG, [cen2021fast]).
Consider natural policy gradient with a tabular softmax policy parametrization, a fixed regularization strength $\tau>0$, and stepsize $0<\eta\leq\frac{1-\gamma}{\tau},$ and denote the iterates of the natural policy gradient updates by $(\pi_{k})_{k\in\mathbb{N}}$. Then for $k\in\mathbb{N}$, it holds that
@@
Through a direct combination of Lemma 6.3, Theorem 4.2, and Theorem 6.4, we obtain (6.3). ∎
Recently, a sharp asymptotic analysis of entropy-regularized natural policy gradient methods has been conducted in [liu2024elementary] showing $R_{\tau}^{\star}-R(\pi_{k})=O((1+\eta\tau)^{-2k})$ compared to the $O((1-\eta\tau)^{k})$ convergence of Theorem 6.4. However, an analogous estimate on the convergence of the policies, as well as a control on the entry times after which the rate holds, is missing, which prevents us from using it in our analysis. By Theorem 4.9 unregularized natural policy gradients achieve $O(k^{\gamma c}e^{-\Delta\eta k})$ convergence. Hence, although Theorem 6.1 improves existing $O(\frac{\log k}{k})$ guarantees, the provided rate is still asymptotically slower compared to unregularized natural policy gradients. When we consider a fixed regularization and step size, $O(\gamma^{k})$-convergence was established when exponentially decreasing the regularization and increasing the step size [li2023homotopic], which can also be achieved by unregularized policy mirror descent, where the rate $O(\gamma^{k})$ is known to be optimal [johnson2024optimal]. For the small stepsize limit $\eta\to 0$, the updates of the regularized natural policy gradient scheme follow the regularized Kakade gradient flow recently studied in [kerimkulov2023fisher]. These results complement the discrete-time analysis and ensure that $D_{\operatorname{K}}(\pi^{\star}_{\tau},\pi_{t})\leq e^{-\tau t}D_{\operatorname{K}}(\pi^{\star}_{\tau},\pi_{0})$, see [kerimkulov2023fisher, Equation (61)]. An analog treatment to the discrete-time case yields an overall estimate of $O(e^{-\sqrt{{2^{-1}\Delta t}}})$ for the regularized flow with strength $\tau=\sqrt{2\Delta t^{-1}}$, compared to the existing $O(t^{-1})$ guarantee in [sethi2024entropy].
@@ 7. Conclusions
In this work, we provide an explicit characterization of the error introduced by entropy regularization in infinite-horizon discounted Markov decision processes with finite state and action spaces. We show that the solutions of the regularized optimization problems satisfy a gradient flow equation with respect to a Riemannian metric common in natural policy gradient methods. We use this to provide upper and lower bounds on the regularization error that decrease exponentially in the inverse regularization strength and match up to polynomial factors with an instance-dependent exponent. Thereby, we characterize the optimal exponential convergence rate for entropy regularization in discounted Markov decision processes. Further, this correspondence allows us to identify the limit of this gradient flow as the generalized maximum entropy optimal policy, thereby characterizing the implicit bias of this gradient flow, which corresponds to a time-continuous version of the natural policy gradient method. We use our improved error estimates to show that for entropy-regularized natural policy gradient methods, the total error decays exponentially in the square root of the number of iterations.
@@ A.1. Performance difference lemma
We recall an expression of the difference between the rewards of two policies in terms of the advantage function and the state-action distribution, and for a proof, we refer to [agarwal2021theory, Lemma 2].
@@ Appendix B Sublinear Convergence of Unregularized Natural Policy Gradient
Here, we generalize the sublinear convergence analysis for unregularized natural policy gradients of [agarwal2021theory, Section 5.3] to non-uniform initial policies $\pi_{0}$ and unnormalized reward $r\in\mathbb{R}^{\mathscr{S}\times\mathscr{A}}$.
@@
First, note that the natural policy gradient iteration is independent of the initial distribution $\mu\in\Delta_{\mathscr{S}}$. Using the performance difference Lemma A.1 as well as the definition of the natural policy gradient step, we estimate the suboptimality for an arbitrary state $s_{0}\in\mathscr{S}$ via
@@ Acknowledgments
The authors acknowledge funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under the project number 442047500 through the Collaborative Research Center Sparsity and Singular Structures (SFB 1481). SC acknowledges funding by the Federal Ministry of Education and Research (BMBF) and the Ministry of Culture and Science of the German State of North Rhine-Westphalia (MKW) under the Excellence Strategy of the Federal Government and the Länder. JM acknowledges funding by the European Union (ERC, FluCo, grant agreement No. 101088488). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or of the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.